# PUSSM: Point Cloud Upsampling as Implicit Statistical Shape Model

### v3

@@ 4.3 Ablation Studies
As a result, training on the MedShapePelvic dataset performed better than training on PU1k. This indicates that PUCRN has indeed learned Pelvic’s shape features. This is very important. However, Grad-PU on the PU1k dataset performed better than MedShapePelvic’s test set. These reasons are multifaceted, and Grad-PU may be better able to learn shape features and knowledge from artificially synthesized PU1k datasets, rather than being adept at learning features from MedShapePelvic obtained from real volume medical images.
Table 3: Pelvic point cloud upsampling results in two different train sets: Network(trainset).
@@
[…]
@@
4.4 Reconstruction Visualizations
The Pelvic1k subdataset 3,4 and 6 have been inputted into SAM-Med3D, and after medical image segmentation, we obtain predicted masks of the pelvic bone, which are then stacked to obtain voxels and converted into point clouds. The current point cloud is sparse, and the segmentation point cloud we obtained is standardized to input 65,536 points into the upsampling model and perform 16×16$\times$16 × upsampling. The model surface is reconstructed using the Voxel Grid with Marching Cubes[20], where voxel size is $1.5$. As shows in Figure 4, compared with the surface of ground truth, and obtained by directly segmenting voxels through marching cubes. Our proposed framework achieves smoother after upsampling in both Grad-PU and PUCRN and more complete shape reconstruction in PUCRN. And there are no obvious contours of layer after reconstructed.
4.5 Reconstruction Evaluations
To evaluate the quality of the mesh surface vis-a-vis ground truth, we used three metrics in Section3.4.2 For Surface Reconstruction.
Table 4: Surface reconstruction results on Pelvic1k subdataset by mesh quality metrics: Area-Length Ratio (ALR), Manifoldness Rate (MR), and Connected Component Discrepancy (CC Diff.).
@@ 4.5 Reconstruction Evaluations
PUSSM (by Grad-PU) 0.811 100.0 2.77 0.817 99.9 9.66 0.819 100.0 7.77
PUSSM (by PUCRN) 0.821 100.0 2.06 0.817 100.0 5.72 0.824 100.0 5.34
As shows in Table 4. Our method outperforms ground truth and segmented result in terms of mesh quality and achieves stable reconstruction accuracy. Specifically, our method has uniform grid partitioning (as shown in ALR) and eliminates floating artifacts (as shown in CC Diff.). This improvement is attributed to the finer points and distribution brought about by point cloud upsampling.
Furthermore, we conducted quantitative comparisons on the reconstruction results in order to evaluate surface reconstruction accuracy and feature preservation. We used metrics (CD, F1, NC, ECD, EF1) in Section3.4.2 For Surface Reconstruction.
Table 5: Surface reconstruction quantitative results on on Pelvic1k subdataset with 16,384 points by poisson sampling.
@@
PUSSM (by Grad-PU) 7.200 0.223 0.671 7.326 0.445 53.297 0.109 0.651 50.104 0.254 12.024 0.223 0.658 8.351 0.453
PUSSM (by PUCRN) 7.111 0.229 0.766 6.295 0.464 51.799 0.118 0.678 49.824 0.266 11.620 0.229 0.686 8.184 0.476
@@
As shows in Table 5. Our method, across all three subdatasets, PUCRN consistently outperforms both Grad-PU and the raw SAM-Med3D segmentations. For example, on Subdataset 3, PUCRN achieves a significantly lower CD compared to Grad-PU, along with superior edge preservation as reflected by a lower ECD and a higher EF1-score. Similar trends are observed on Subdataset 4 and Subdataset 6, where PUCRN achieves the best overall performance in terms of both global surface quality and local edge fidelity. These results demonstrate the effectiveness of PUCRN in enhancing the downstream surface reconstruction quality by providing denser and more accurate point distributions.
5 Conclusion
@@
In this work, we proposed PUSSM, an implicit statistical shape modeling framework that leverages deep point cloud upsampling to reconstruct high-resolution pelvic structures from sparse medical data. Our method avoids the limitations of classical SSMs by eliminating the need for landmark annotations and explicitly modeling shape variation. Experiments demonstrate improved surface quality and anatomical plausibility across multiple evaluation metrics.
Future work will extend this framework to 3D SAM model[15] by points based prompting to integrate it as Chain-of-Thought (CoT)[12] for Segmentation.
{credits}

### v4

@@ IV-A3 Implementation Details
TABLE III: Pelvic point cloud upsampling results in two different train sets: Network(trainset).
@@
$downarrow CD\downarrow HD\downarrow$ $downarrow CD\downarrow HD\downarrow$
@@
IV-B Upsampling Results
In accuracy analysis, as mentioned above, the generalization performance of two SOTA network have been evaluated on the MedShapePelvic dataset in $4\times$ upsampling. The quantitative results for the dataset in Table I show that our method performs consistently across different input scales (512, 1024, 2,048 and 4,096 points) and outperforms the original network in CD, HD, and P2F metrics.
Besides, we also conducted $16\times$ upsampling tests, in Table II. Grad-PU can perform upsampling at any scale, so we set its ratio directly to $16\times$ on the test set. However, the testing model of PUCRN can only use the same upsampling factor as the training conditions, so we performed two $4\times$ samplings on PUCRN to achieve a $16\times$ sampling effect. The results show that PUCRN still has good performance.
IV-C Ablation Studies
To verify the effectiveness and necessity of training in MedShapePelvic to learn the prior knowledge of shape, the $4\times$ sampling result based on training in MedShapePelvic and PU1k dataset[21] have been tested. Among them, MedShapePelvic has 50496 training samples and PU1k has 69000 training samples. This avoids the difference in learning effectiveness caused by the large difference in the number of training samples. The result shows in Table III
@@ IV-C Ablation Studies
As a result, training on the MedShapePelvic dataset performed better than training on PU1k. This indicates that PUCRN has indeed learned Pelvic’s shape features, which means that it learned statistical variation embedded in anatomical structures, serving as an implicit anatomical prior without relying on PCA or landmarks. However, Grad-PU on the PU1k dataset performed better than MedShapePelvic’s test set. These reasons are multifaceted, and Grad-PU may be better able to learn shape features and knowledge from artificially synthesized PU1k datasets, rather than being adept at learning features from MedShapePelvic obtained from real volume medical images.
TABLE IV: Surface reconstruction results on Pelvic1k subdataset by mesh quality metrics: Area-Length Ratio (ALR), Manifoldness Rate (MR), and Connected Component Discrepancy (CC Diff.).
@@
Med-PU (by Grad-PU) 0.811 100.0 2.77 0.817 99.9 9.66 0.819 100.0 7.77
Med-PU (by PUCRN) 0.821 100.0 2.06 0.817 100.0 5.72 0.824 100.0 5.34
IV-D Reconstruction Visualizations
TABLE V: Surface reconstruction quantitative results on on Pelvic1k subdataset with 16,384 points by poisson sampling.
@@ IV-D Reconstruction Visualizations
Med-PU (by Grad-PU) 7.200 0.223 0.671 7.326 0.445 53.297 0.109 0.651 50.104 0.254 12.024 0.223 0.658 8.351 0.453
Med-PU (by PUCRN) 7.111 0.229 0.766 6.295 0.464 51.799 0.118 0.678 49.824 0.266 11.620 0.229 0.686 8.184 0.476
The Pelvic1k subdataset 3,4 and 6 have been inputted into SAM-Med3D, and after medical image segmentation, we obtain predicted masks of the pelvic bone, which are then stacked to obtain voxels and converted into point clouds. The current point cloud is sparse, and the segmentation point cloud we obtained is standardized to input 65,536 points into the upsampling model and perform $16\times$ upsampling. The model surface is reconstructed using the Voxel Grid with Marching Cubes[43], where voxel size is $1.5$. As shows in Figure 3, compared with the surface of ground truth, and obtained by directly segmenting voxels through marching cubes. Our proposed framework achieves smoother after upsampling in both Grad-PU and PUCRN and more complete shape reconstruction in PUCRN. And there are no obvious contours of layer after reconstructed.
IV-E Reconstruction Evaluations
To evaluate the quality of the mesh surface vis-a-vis ground truth, we used three metrics in SectionIII-C2 For Surface Reconstruction.
As shows in Table IV. Our method outperforms ground truth and segmented result in terms of mesh quality and achieves stable reconstruction accuracy. Specifically, our method has uniform grid partitioning (as shown in ALR) and eliminates floating artifacts (as shown in CC Diff.). This improvement is attributed to the finer points and distribution brought about by point cloud upsampling.
Furthermore, we conducted quantitative comparisons on the reconstruction results in order to evaluate surface reconstruction accuracy and feature preservation. We used metrics (CD, F1, NC, ECD, EF1) in SectionIII-C2 For Surface Reconstruction.
@@ IV-E Reconstruction Evaluations
As shows in Table V. Our method, across all three subdatasets, PUCRN consistently outperforms both Grad-PU and the raw SAM-Med3D segmentations. For example, on Subdataset 3, PUCRN achieves a significantly lower CD compared to Grad-PU, along with superior edge preservation as reflected by a lower ECD and a higher EF1-score. Similar trends are observed on Subdataset 4 and Subdataset 6, where PUCRN achieves the best overall performance in terms of both global surface quality and local edge fidelity. These results demonstrate the effectiveness of PUCRN in enhancing the downstream surface reconstruction quality by providing denser and more accurate point distributions. And we observe consistent reductions in Chamfer Distance and improvements in Normal Consistency and F-scores, indicating that the upsampling network effectively maps sparse point clouds toward the anatomical shape manifold as Figure 2.
V Discussion
This study demonstrates that coupling a robust volumetric segmenter with a dedicated point cloud upsampler can bridge the gap between voxel labels and mesh-quality anatomical models. Compared with classical SSMs that rely on landmarks and explicit statistical spaces, our approach learns an implicit prior from realistic pelvic shapes (MedShapePelvic), enabling automatic densification and regularization of sparse, segmentation-derived points into smooth, watertight surfaces suited for visualization and downstream analysis.
Why domain supervision matters
Our ablations indicate consistent gains when the upsampler is trained on anatomical meshes (MedShapePelvic) rather than synthetic/CAD-style objects (PU1k). This aligns with the intuition that the statistics of cortical thickness, local curvature, and trabecular recesses in pelvic bones differ markedly from generic object datasets. Prior CMPB studies similarly show that medical-domain supervision/priors improve realism and downstream usability of the reconstructed geometry [44, 45]. In practice, domain-aligned supervision helps suppress perforations and over-smoothing, yielding lower HD and improved edge fidelity after meshing.
Limitations and Future Work
The performance of our 3D reconstruction pipeline, Med-PU, is limited by several factors. First, it depends on the quality of front-end segmentation and is sensitive to imaging domain shifts, such as challenging boundaries or metal artifacts, which may propagate errors to the point set and surface. Second, our meshing relies on iso-surface extraction without explicit topology control, and while we apply standard cleanup and report manifoldness-oriented metrics, pathological non-manifold cases may still occur. Third, our evaluation is geometry-centric, and the clinical impact—such as interobserver variability in contouring, planning accuracy, or dose/navigation surrogates—remains to be quantified through future user studies. Broader evaluations across diverse anatomies, scanners, and acquisition protocols are needed to assess robustness to domain shifts. To address inherited segmentation errors, joint or end-to-end training of segmentation and upsampling could be explored. Additionally, runtime and memory demands increase with higher upsampling ratios, suggesting the need for lighter backbones, GPU-friendly operators, or uncertainty-aware pruning to enhance efficiency and reliability. Finally, integrating task-level endpoints and extending the implicit prior beyond medical applications (e.g., to robotics or industrial inspection) would further validate Med-PU as a general-purpose, knowledge-based 3D reconstruction system. Broader surveys, such as those in clinical 3D reconstruction for endoscopy, emphasize the importance of task-driven validation beyond geometric scores [46].
VI Conclusion
@@
In this work, we proposed Med-PU, an implicit anatomical prioring framework that leverages deep point cloud upsampling to reconstruct high-resolution pelvic structures from sparse medical data. Our method avoids the limitations of classical SSMs by eliminating the need for landmark annotations and explicitly modeling shape variation. Experiments demonstrate improved surface quality and anatomical plausibility across multiple evaluation metrics.
Instead of explicitly decomposing shape variations, Med-PU learns to embed these variations in a function space. This allows the model to restore anatomically plausible shapes from sparse or noisy segmentations, demonstrating its ability to capture and regularize shape diversity.